Wendy is resting on her bike at the top of a 225 meter tall hill. She releases the brake and begins to roll down the hill. If Wendy and the bike has a total mass of 75.0 kg, what is the speed of the bike when it reaches the bottom of the hill and begins coasting?
1
Expert's answer
2020-11-17T13:03:59-0500
"v=\sqrt{2gh}\\v=\sqrt{2(9.8)(225)}=66.4\frac{m}{s}"